# Neyman-Pearson multiclass classification under label noise via empirical likelihood

### 23 Mar 2026 v1

@@ 3 Empirical likelihood based estimation under label noise
Since we observe contaminated labels $\widetilde{Y}\in[K]$ rather than the true labels $Y$ in the noisy label setting. Conventional approaches that ignore this label noise yield estimators for $\{\widetilde{w}_{k}^{*},\widetilde{\pi}_{k}^{*}(x)\}_{k=0}^{K-1}$, which differ from the target quantities $\{w_{k}^{*},\pi_{k}^{*}(x)\}_{k=0}^{K-1}$ needed for proper NP type error control. To address this, we propose a flexible approach that directly estimates $\{w_{k}^{*},\pi_{k}^{*}(x)\}_{k=0}^{K-1}$ from noisy labeled data. Our approach uses density ratio modeling and empirical likelihood, without requiring any prior knowledge of the label corruption probabilities.
This section presents our estimator along with its numerical computation and statistical properties for general classification problems with label noise.
@@ 3.1 Exponential tilting and density ratio model
$\frac{dP_{k}^{*}}{dP_{0}^{*}}(x)=\exp(\gamma_{k}^{*}+\langle\beta_{k}^{*},g(x)\rangle),$ (7)
@@
$\int\exp(\gamma_{k}^{*}+\langle\beta_{k}^{*},g(x)\rangle)P_{0}^{*}(dx)=1,\quad\forall~k\in 1,\ldots,K-1.$ (8)
@@
Both (7) and (9) are instances of the density ratio model (DRM) (Anderson, 1979; Qin, 2017), where all measures ($\{\widetilde{P}_{l}^{*}\}_{l=0}^{K-1}$, $\{P_{k}^{*}\}_{k=1}^{K-1}$) are connected to the base measure $P_{0}^{*}$ via a parametric form. The DRM is highly flexible and encompasses many commonly used distribution families, including all members of the exponential family with appropriate specifications for $P_{0}^{*}$ and $g(x)$. For example, if the base distribution is normal and $g(x)=(x,x^{2})^{\top}$, the DRM reduces to the normal distribution family.
@@
The function $g$ is commonly referred to as the basis function. In our simulations and experiments, the simple choice $g(x)=x$ performs well across a range of scenarios. For more complex data, such as images, performance may be improved by using features extracted from pretrained deep neural networks or by adopting data-adaptive approaches, such as that of Zhang and Chen (2022).
@@ 3.2 Identifiability
In this section, we study the identifiability of the model in (9). We emphasize that, in the absence of additional structural assumptions, the model is fundamentally unidentifiable—beyond the usual label permutation ambiguity encountered in mixture models. In particular, multiple distinct parameterizations can induce exactly the same observed distributions. We illustrate this issue through two examples below.
Example 3.1 (Algebraic unidentifiability of true distributions).
If $P_{0}^{*}(x)$ and $P_{1}^{*}(x)$ are unconstrained, the model is unidentifiable. For any small constant $\epsilon\in(0,1)$, we can define a new, "contaminated" candidate distribution for class 1:
$P_{1}^{**}(x)\triangleq(1-\epsilon)P_{1}^{*}(x)+\epsilon P_{0}^{*}(x).$
Because $P_{1}^{**}(x)$ is a convex combination of two valid probability density functions, it is also a valid density function. We can rewrite the true $P_{1}^{*}(x)$ in terms of this new distribution:
$\widetilde{P}_{0}(x)$ $=(1-\rho_{0}^{*})P_{0}^{*}(x)+\rho_{1}^{*}\left[\frac{1}{1-\epsilon}P_{1}^{**}(x)-\frac{\epsilon}{1-\epsilon}P_{0}^{*}(x)\right],$
$=\left(1-\rho_{0}^{*}-\frac{\epsilon\rho_{1}^{*}}{1-\epsilon}\right)P_{0}^{*}(x)+\left(\frac{\rho_{1}^{*}}{1-\epsilon}\right)P_{1}^{**}(x).$
@@
Recall our density ratio model formulation from (7):
$\frac{dP_{1}^{*}}{dP_{0}^{*}}(x)=\exp(\gamma_{1}^{*}+\langle\beta_{1}^{*},g(x)\rangle).$
Suppose the chosen feature representation $g(x)=(g_{1}(x),g_{2}(x))^{\top}\in{\mathbb{R}}^{2}$ is not linearly independent across the support of $P_{0}^{*}$. Specifically, assume the data strictly lies on a manifold where $g_{1}(x)+g_{2}(x)=c$ for some constant $c$.
Let $(\gamma_{1}^{*},\beta_{11}^{*},\beta_{12}^{*})$ be the true parameters generating the data. The exponential tilt is:
$\gamma_{1}^{*}+\beta_{11}^{*}g_{1}(x)+\beta_{12}^{*}g_{2}(x).$
Now, consider a different set of parameters defined by shifting the feature coefficients by an arbitrary constant $\delta\neq 0$:
$\beta_{11}^{**}=\beta_{11}^{*}+\delta,\quad\beta_{12}^{**}=\beta_{12}^{*}+\delta.$
Evaluating the tilt with these new parameters yields:
$\gamma_{1}^{*}+\beta_{11}^{**}g_{1}(x)+\beta_{12}^{**}g_{2}(x)$ $=\gamma_{1}^{*}+(\beta_{11}^{*}+\delta)g_{1}(x)+(\beta_{12}^{*}+\delta)g_{2}(x)$
$=\gamma_{1}^{*}+\beta_{11}^{*}g_{1}(x)+\beta_{12}^{*}g_{2}(x)+\delta(g_{1}(x)+g_{2}(x))$
$=(\gamma_{1}^{*}+\delta c)+\beta_{11}^{*}g_{1}(x)+\beta_{12}^{*}g_{2}(x).$
By defining a new baseline parameter $\gamma_{1}^{**}=\gamma_{1}^{*}+\delta c$, we perfectly absorb the shift. The parameter sets $(\gamma_{1}^{*},\beta_{11}^{*},\beta_{12}^{*})$ and $(\gamma_{1}^{**},\beta_{11}^{**},\beta_{12}^{**})$ produce the exact same density ratio $\exp(\gamma_{1}^{**}+\langle\beta_{1}^{**},g(x)\rangle)\equiv\exp(\gamma_{1}^{*}+\langle\beta_{1}^{*},g(x)\rangle)$ for all observable $x$. Thus, without assumptions guaranteeing the richness of the feature subspace (i.e., ruling out collinearity), the exponential tilt parameters are non-identifiable.
These examples highlight that the model is intrinsically unidentifiable without further structural constraints. To resolve this issue, we impose a set of mild but essential assumptions that ensure identifiability, as stated in the following theorem.
@@
[…]
@@ 3.3 Empirical likelihood based inference
The log-likelihood function based on the contaminated set ${\mathcal{D}}=\{(X_{i},\widetilde{Y}_{i})\}_{i=1}^{n}\overset{\text{i.i.d.}}{\sim}P^{*}_{X,\widetilde{Y}}$ is
$\ell=\sum_{i=1}^{n}\sum_{l=0}^{K-1}\left[\mathbbm{1}(\widetilde{Y}_{i}=l)\log\left\{\widetilde{w}_{l}^{*}\widetilde{P}_{l}^{*}(\{X_{i}\})\right\}\right].$
Using the exponential tilting relationship in (9), we obtain
@@
The empirical likelihood (EL) (Owen, 2001; Qin and Lawless, 1994) allows us to estimate parameters without assumptions on the form of $P_{0}^{*}$. Specifically, let $p_{i}=P_{0}^{*}(\{X_{i}\})$ for $i=1,\ldots,n$ and view them as parameters. Let ${\bm{w}}=\{w_{k}\}_{k=0}^{K-1}$ and $\mbox{$\theta$}=\{{\bm{w}},\mbox{$\gamma$},\mbox{$\beta$},{\bm{T}}\}$. The log-EL function becomes:
@@
[…]
@@
$\sum_{i=1}^{n}\frac{\exp(\gamma_{k}+\beta_{k}^{\top}g(X_{i}))-1}{1+\sum_{k^{\prime}=1}^{K-1}\nu_{k^{\prime}}\left\{\exp(\gamma_{k^{\prime}}+\beta_{k^{\prime}}^{\top}g(X_{i}))-1\right\}}=0,\quad k=1,\ldots,K-1.$ (13)
@@
$\widehat{P}_{k}=\sum_{i=1}^{n}p_{i}(\widehat{}\mbox{$\theta$})\exp(\widehat{\gamma}_{k}+\widehat{\beta}_{k}^{\top}g(X_{i}))\delta_{X_{i}},\quad\widehat{\pi}_{k}(x)=\frac{\exp(\widehat{\gamma}_{k}^{\dagger}+\widehat{\beta}_{l}^{\top}g(x))}{\sum_{k^{\prime}}\exp(\widehat{\gamma}_{k^{\prime}}^{\dagger}+\widehat{\beta}_{k^{\prime}}^{\top}g(x))},$ (15)
@@ 3.4 Statistical guarantees
Let $\mbox{$\theta$}^{*}=({\bm{w}}^{*},\mbox{$\gamma$}^{*},\mbox{$\beta$}^{*},{\bm{T}}^{*})$ be the true values of $\theta$ and $\widehat{}\mbox{$\theta$}$ be its MELE estimator in (14). We have
$\sqrt{n}(\widehat{}\mbox{$\theta$}-\mbox{$\theta$}^{*})\to N(0,\Sigma)$
as $n\to\infty$ where $\Sigma>0$ is some non-degenerated covariance matrix.
[…]
@@ 3.5 EM algorithm for profile EL maximization
1:Noisy labeled data $\{(X_{i},\widetilde{Y}_{i})\}_{i=1}^{n}$, feature mapping $g(x)$, number of classes $K$, convergence threshold $\epsilon$
@@
Using the chain rule and the instance-independent Assumption 2.1, we decompose the joint probability into:
$\begin{split}{\mathbb{P}}(X=X_{i},\widetilde{Y}_{i}=l,Y_{i}=k)&={\mathbb{P}}(X=X_{i}|\widetilde{Y}_{i}=l,Y_{i}=k){\mathbb{P}}(\widetilde{Y}_{i}=l|Y_{i}=k){\mathbb{P}}(Y_{i}=k)\\ &={\mathbb{P}}(X=X_{i}|Y_{i}=k)T_{lk}^{*}w_{k}^{*}\\ &=P_{0}^{*}(\{X_{i}\})\exp\{\gamma_{k}^{*}+\langle\beta_{k}^{*},g(X_{i})\rangle\}T_{lk}^{*}w_{k}^{*},\end{split}$
@@
We define the profile complete data log-EL as
$p\ell_{n}^{c}(\mbox{$\theta$})=\sup_{{\bm{p}}}\ell_{n}^{c}({\bm{p}},\mbox{$\theta$})$
where the maximization is subject to the constraints in (11). Using the same derivation for (12), the profile complete data log-EL becomes
@@
  • E-step (Expectation step): The missing data is “filled in” by computing an educated guess based on the current parameter estimate $\mbox{$\theta$}^{(t)}$. This involves calculating the posterior distribution of the missing data given the observed data and the current parameters:
    $\begin{split}\omega_{ik}^{(t)}=&~{\mathbb{E}}(\mathbbm{1}(Y_{i}=k)|\mbox{$\theta$}^{(t)};\widetilde{Y}_{i},X_{i})={\mathbb{P}}(Y_{i}=k|\mbox{$\theta$}^{(t)};\widetilde{Y}_{i},X_{i})\\ =&~\frac{\exp(\gamma_{k}^{(t)}+\langle\beta_{k}^{(t)},g(X_{i})\rangle)T_{\widetilde{Y}_{i},k}^{(t)}w_{k}^{(t)}}{\sum_{k^{\prime}=1}^{K}\exp(\gamma_{k^{\prime}}^{(t)}+\langle\beta_{k^{\prime}}^{(t)},g(X_{i})\rangle)T_{\widetilde{Y}_{i},k^{\prime}}^{(t)}w_{k^{\prime}}^{(t)}}.\end{split}$ (16)
@@
[…]
@@
    The update of $(\mbox{$\gamma$}^{(t+1)},\mbox{$\beta$}^{(t+1)})$ is slightly more complicated. Note that $(\mbox{$\gamma$}^{(t+1)},\mbox{$\beta$}^{(t+1)})$ maximizes the function
@@
[…]
@@
[…]
@@
The E-step and M-step are repeated iteratively until the change in the profile log-EL falls below a predefined threshold. This stopping criterion is guaranteed to be met because the EM algorithm produces a sequence of estimates that monotonically increases the true objective—the profile log-EL, which we detail below.
@@
[…]
@@
[…]

### 21 Apr 2026 v2

@@ 3 Empirical likelihood based estimation under label noise
Motivated by the need for structural assumptions identified in the previous section, we develop an estimation framework that directly recovers $\{w_{k}^{*},\pi_{k}^{*}(x)\}_{k=0}^{K-1}$ from noisy labeled data. Our approach combines density ratio modeling with empirical likelihood and does not require prior knowledge of the label corruption probabilities.
@@ 3.1 Exponential tilting and density ratio model
$dP_{k}^{*}/dP_{0}^{*}(x)=\exp(\gamma_{k}^{*}+\langle\beta_{k}^{*},g(x)\rangle),$ (6)
@@
$\int\exp(\gamma_{k}^{*}+\langle\beta_{k}^{*},g(x)\rangle)P_{0}^{*}(dx)=1,\qquad\forall~k\in\{1,\ldots,K-1\}.$ (7)
@@
Both (6) and (8) are instances of the density ratio model (DRM) (Anderson, , 1979; Qin, , 2017), where all measures ($\{\widetilde{P}_{l}^{*}\}_{l=0}^{K-1}$, $\{P_{k}^{*}\}_{k=1}^{K-1}$) are connected to the base measure $P_{0}^{*}$ via a parametric form. The DRM is highly flexible and encompasses many commonly used distribution families, including all members of the exponential family with appropriate specifications for $P_{0}^{*}$ and $g(x)$. For example, if the base distribution is normal and $g(x)=(x,x^{2})^{\top}$, the DRM reduces to the normal distribution family.
@@
The function $g$ is commonly referred to as the basis function. In our simulations and experiments, the simple choice $g(x)=x$ performs well across a range of scenarios. For more complex data, such as images, performance may be improved by using features extracted from pretrained deep neural networks or by adopting data-adaptive approaches, such as that of Zhang and Chen, (2022).
@@ 3.2 Identifiability
In this section, we study the identifiability of the model in (8). We first show that the DRM is not identifiable when the features are linearly dependent, highlighting that the absence of collinearity is essential for establishing identifiability of the model in (8).
@@
Recall our density ratio model formulation from (6): $dP_{1}^{*}/dP_{0}^{*}(x)=\exp(\gamma_{1}^{*}+\langle\beta_{1}^{*},g(x)\rangle)$. Suppose the chosen feature representation $g(x)=(g_{1}(x),g_{2}(x))^{\top}\in{\mathbb{R}}^{2}$ is not linearly independent across the support of $P_{0}^{*}$. Specifically, assume the data lies on a manifold where $g_{1}(x)+g_{2}(x)=c$ for some constant $c$.
Let $(\gamma_{1}^{*},\beta_{11}^{*},\beta_{12}^{*})$ be the true parameters generating the data. The exponential tilt is: $\gamma_{1}^{*}+\beta_{11}^{*}g_{1}(x)+\beta_{12}^{*}g_{2}(x)$. Now, consider a different set of parameters defined by shifting the feature coefficients by an arbitrary constant $\delta\neq 0$: $\beta_{11}^{**}=\beta_{11}^{*}+\delta$, $\beta_{12}^{**}=\beta_{12}^{*}+\delta.$ Evaluating the tilt with these new parameters yields:
$\begin{split}\gamma_{1}^{*}+\beta_{11}^{**}g_{1}(x)+\beta_{12}^{**}g_{2}(x)&=\gamma_{1}^{*}+(\beta_{11}^{*}+\delta)g_{1}(x)+(\beta_{12}^{*}+\delta)g_{2}(x)\\ &=(\gamma_{1}^{*}+\delta c)+\beta_{11}^{*}g_{1}(x)+\beta_{12}^{*}g_{2}(x).\end{split}$
By defining a new baseline parameter $\gamma_{1}^{**}=\gamma_{1}^{*}+\delta c$, we perfectly absorb the shift. The parameter sets $(\gamma_{1}^{*},\beta_{11}^{*},\beta_{12}^{*})$ and $(\gamma_{1}^{**},\beta_{11}^{**},\beta_{12}^{**})$ produce the exact same density ratio $\exp(\gamma_{1}^{**}+\langle\beta_{1}^{**},g(x)\rangle)\equiv\exp(\gamma_{1}^{*}+\langle\beta_{1}^{*},g(x)\rangle)$ for all observable $x$. Thus, without assumptions guaranteeing the richness of the feature subspace (i.e., ruling out collinearity), the tilt parameters are non-identifiable.
This example motivates the introduction of a set of mild but essential assumptions under which identifiability can be restored, as formalized in the following theorem.
@@
Under mild assumptions (Assumptions C.1– C.3), the parameters $({\bm{M}}^{*},\mbox{$\gamma$}^{*},\mbox{$\beta$}^{*})$ and the reference distribution $P_{0}^{*}$ in (8) are strictly identifiable.
The formal mathematical statements of these assumptions, along with the complete proof, are deferred to Appendix C.4. Strict identifiability implies a true one-to-one mapping between the parameter space and the probability distributions of the observed data. For our model, this means there is exactly one unique set of parameters $({\bm{M}}^{*},\mbox{$\gamma$}^{*},\mbox{$\beta$}^{*})$ and one unique reference distribution $P_{0}^{*}$ that can generate the observed data, completely eliminating the label-switching (permutation) ambiguity that typically plagues mixture models. To establish this strict identifiability, we require three standard regularity conditions. Broadly, these conditions ensure that the latent mixture components are physically distinct and that the observed data provides sufficient variation to decouple them. Specifically, we require the feature coefficients of the exponential tilts to be strictly distinct (Assumption C.1) and the feature mapping to explore a sufficiently rich subspace (Assumption C.2). These are common requirements in the exponential family and finite mixture model literature to prevent component collapse and guarantee that the underlying densities are linearly independent. Furthermore, we require the confusion matrix $\bm{M}$$^{*}$ to be full rank, with specific structural constraints on its inverse (Assumption C.3). Crucially, this final condition on the inverse matrix not only ensures the signal from the true latent components is preserved, but it effectively breaks the permutation symmetry inherent in standard mixture models, allowing for the strict, unique identification of all parameters.
@@ 3.3 Empirical likelihood based inference
The log-likelihood function based on the contaminated set ${\mathcal{D}}=\{(X_{i},\widetilde{Y}_{i})\}_{i=1}^{n}\overset{\text{i.i.d.}}{\sim}P^{*}_{X,\widetilde{Y}}$ is $\ell=\sum_{i=1}^{n}\sum_{l=0}^{K-1}\left[\mathbbm{1}(\widetilde{Y}_{i}=l)\log\left\{\widetilde{w}_{l}^{*}\widetilde{P}_{l}^{*}(\{X_{i}\})\right\}\right]$. Using the exponential tilting relationship in (8), we obtain
@@
[…]
@@
The inference for $\theta$ is usually made by first profiling the log-EL with respect to $\bm{p}$}. That is, we define $p\ell_{n}(\mbox{$\theta$})=\sup_{{\bm{p}}}\ell_{n}({\bm{p}},\mbox{$\theta$})$ subject to the constraints in (10). By the Lagrange multiplier method (see Appendix C.2 for details), the maximizer is attained when
@@
$\sum_{i=1}^{n}\frac{\exp(\gamma_{k}+\beta_{k}^{\top}g(X_{i}))-1}{1+\sum_{k^{\prime}=1}^{K-1}\nu_{k^{\prime}}\left\{\exp(\gamma_{k^{\prime}}+\beta_{k^{\prime}}^{\top}g(X_{i}))-1\right\}}=0,\qquad k=1,\ldots,K-1.$ (12)
@@
$\widehat{P}_{k}=\sum_{i=1}^{n}p_{i}(\widehat{}\mbox{$\theta$})\exp(\widehat{\gamma}_{k}+\widehat{\beta}_{k}^{\top}g(X_{i}))\delta_{X_{i}},\qquad\widehat{\pi}_{k}(x)=\frac{\exp(\widehat{\gamma}_{k}^{\dagger}+\widehat{\beta}_{l}^{\top}g(x))}{\sum_{k^{\prime}}\exp(\widehat{\gamma}_{k^{\prime}}^{\dagger}+\widehat{\beta}_{k^{\prime}}^{\top}g(x))},$ (14)
@@ 3.4 Statistical guarantees
Let $\mbox{$\theta$}^{*}=({\bm{w}}^{*},\mbox{$\gamma$}^{*},\mbox{$\beta$}^{*},{\bm{T}}^{*})$ be the true values of $\theta$ and $\widehat{}\mbox{$\theta$}$ be its MELE estimator in (13). We have $\sqrt{n}(\widehat{}\mbox{$\theta$}-\mbox{$\theta$}^{*})\to N(0,\Sigma)$ as $n\to\infty$ where $\Sigma>0$ is some non-degenerated covariance matrix.
The proof of the theorem is deferred to Appendix C.3. The result establishes that the MELE estimator is consistent and asymptotically normal, achieving the optimal convergence rate. This asymptotic normality plays a key role in deriving the NP oracle inequalities for the final classifier.
@@ 3.5 EM algorithm for profile EL maximization
1:Noisy labeled data $\{(X_{i},\widetilde{Y}_{i})\}_{i=1}^{n}$, basis $g(x)$, # of classes $K$, convergence threshold $\epsilon$
@@
Using the chain rule and Assumption 2.1, we decompose the joint probability as:
${\mathbb{P}}(X=X_{i},\widetilde{Y}_{i}=l,Y_{i}=k)={\mathbb{P}}(X=X_{i}|Y_{i}=k)T_{lk}^{*}w_{k}^{*}=P_{0}^{*}(\{X_{i}\})\exp\{\gamma_{k}^{*}+\langle\beta_{k}^{*},g(X_{i})\rangle\}T_{lk}^{*}w_{k}^{*},$
@@
We define the profile complete data log-EL as $p\ell_{n}^{c}(\mbox{$\theta$})=\sup_{{\bm{p}}}\ell_{n}^{c}({\bm{p}},\mbox{$\theta$})$, where the maximization is subject to the constraints in (10). Using the same derivation for (11), the profile complete data log-EL becomes
@@
The EM algorithm is an iterative procedure initialized at $\mbox{$\theta$}^{(0)}$. At iteration $t$, given the current estimate $\mbox{$\theta$}^{(t)}$, it proceeds in two steps:
  • E-step (Expectation step): The missing data is “filled in” by computing their posterior given the observed data and $\mbox{$\theta$}^{(t)}$:
    $\omega_{ik}^{(t)}={\mathbb{E}}(\mathbbm{1}(Y_{i}=k)|\mbox{$\theta$}^{(t)};\widetilde{Y}_{i},X_{i})=\frac{\exp(\gamma_{k}^{(t)}+\langle\beta_{k}^{(t)},g(X_{i})\rangle)T_{\widetilde{Y}_{i},k}^{(t)}w_{k}^{(t)}}{\sum_{k^{\prime}=1}^{K}\exp(\gamma_{k^{\prime}}^{(t)}+\langle\beta_{k^{\prime}}^{(t)},g(X_{i})\rangle)T_{\widetilde{Y}_{i},k^{\prime}}^{(t)}w_{k^{\prime}}^{(t)}}.$ (15)
@@
  • M-step (Maximization step): Instead of maximizing log-EL, the M-step maximizes $Q$ with respect to $\theta$ instead, i.e., $\mbox{$\theta$}^{(t+1)}=\operatorname*{arg\,max}Q(\mbox{$\theta$},\mbox{$\theta$}^{(t)})$.
    Since $Q$ is separable in $\bm{w}$}, $\bm{T}$}, and $(\mbox{$\gamma$},\mbox{$\beta$})$, we can optimize each set of parameters separately. Below are the updated parameters; the detailed derivations are provided in Appendix A.1.
@@
    The update of $(\mbox{$\gamma$}^{(t+1)},\mbox{$\beta$}^{(t+1)})$ is more involved, as it maximizes
@@
    At the optimal point, we show in Appendix A.1 that the Lagrange multipliers
@@
    which is the weighted log-likelihood of multinomial logistic regression model based on the dataset in Table 4 in Appendix A.1. Thus, $(\bar{\mbox{$\gamma$}}^{(t+1)},\mbox{$\beta$}^{(t+1)})$ maximizes the weighted log-likelihood. We use standar packages in R or Python to find the numerical value. Finally, we get
@@
The E- and M-steps are iterated until the change in the profile log-EL falls below a threshold. This criterion is met since the EM algorithm produces estimates that monotonically increase the true objective—the profile log-EL, which we detail below.
@@
The proof of Proposition 3.1 is in Appendix A.2. Note that $p\ell_{n}(\mbox{$\theta$}^{(t+1)})$ is a sum of the logarithm of probabilities, implying that $p\ell_{n}(\mbox{$\theta$})\leq 0$ for any $\theta$ that is feasible. Proposition 3.1 then guarantees convergence to at least a local maximum for the given initial value $\mbox{$\theta$}^{(0)}$. We recommend using multiple initial values to explore the likelihood function to ensure that the algorithm reaches the global maximum. Second, in practice, we may stop the algorithm when the increment in the log-EL after an iteration is no greater than, say, $10^{-6}$.
@@
When prior knowledge about the transition matrix is available (e.g., $T_{kk}^{*}\geq\xi_{k}>0$), the EM algorithm’s M-step can be modified to incorporate such prior information. The analytical solution for this constrained update is provided in A.3. To further improve numerical stability and convergence speed, penalties can be imposed on the diagonal elements $\{T_{kk}\}_{k=0}^{K-1}$. The corresponding analytical solution for this penalized update is detailed in Appendix A.4.